OpenStudy (anonymous):

Subtracting 766 from 941 results in a difference of

OpenStudy (saifoo.khan):

175

myininaya (myininaya):

941 -766 ------ 1 7 5

OpenStudy (anonymous):

hey use calculator

myininaya (myininaya):

11-6=5 13-6=7 8-7=1